OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 2) and a directrix at y = -2.

OpenStudy (mathmale):

p represents the distance from the vertex of y our parabola to the focus. p also represents the distance of the directrix from the vertex.

OpenStudy (mathmale):

Decide: is this a vertical parabola or a horiz. one?

OpenStudy (anonymous):

vertical.

OpenStudy (mathmale):

Yes. The standard equation for a vertical parabola involving the parameter p is 4py=x^2. Since p is known (it's 2), replace p with 2 in the above eq'n . what do you get?

OpenStudy (anonymous):

8y=x^2

OpenStudy (mathmale):

yes, and now divide both sides of this equation by 8. Your results?

OpenStudy (anonymous):

y=1/8x^2

OpenStudy (mathmale):

Yes, except that you need to enclose that 1/8 in parentheses for clarity. y=(1/8)x^2. Done. Very nice.

OpenStudy (anonymous):

Thanks that was so much easier than I thought it would be!

OpenStudy (mathmale):

Write down, study and remember this formula for a vertical parabola: 4py=x^2.

OpenStudy (mathmale):

Vertex form: y-k = a(x-h)^2 Standard form: ax^2 + bx + c = 0. Happy to work with you.

OpenStudy (anonymous):

can you help me with another one that is slightly different but the same concept?

OpenStudy (mathmale):

Yes, but I'm afraid my involvement will have to take place later this evening. I'd suggest you post your question now; there are others who understand these matters. Tag me if you need help after trying that.

OpenStudy (anonymous):

Okay thank you.